# Reconstruction Alignment Improves Unified Multimodal Models

### v1

@@ Appendix D More Quantitative Results on Text-to-Image Generation
Why the improvements on Show-o is modest? (I) The CLIP vision encoder exhibits insufficient capacity for extracting high-level semantic information necessary for enhanced generation. (II) Show-o’s constrained codebook size of 4096 tokens fundamentally limits its representational capacity and generation flexibility. (III) Show-o demonstrates minimal inherent image reconstruction capabilities, producing reconstructions with significant noise artifacts. Consequently, RecA must simultaneously learn to mitigate noise generation while conditioning on CLIP features, presenting additional optimization challenges.
Why the text-to-image generation improvements on BAGEL is modest? (I) The computational constraints associated with training the 14B parameter model prevent full exploitation of RecA’s potential benefits. (II) We hypothesize that BAGEL’s pre-existing image editing training, which conditions on both SigLIP and VAE features, has already endowed the model with robust capabilities for SigLIP feature-based reconstruction. This pre-established competency constrains the potential improvement space for RecA in generation tasks, while correspondingly enabling substantial enhancements in image editing capabilities.

### v2

@@ Appendix C Discussion
Why the improvements on Show-o is modest? (I) The CLIP vision encoder exhibits insufficient capacity for extracting high-level semantic information necessary for enhanced generation (Tong et al., 2024b). (II) Show-o’s constrained codebook size of 4096 tokens fundamentally limits its representational capacity and generation flexibility. (III) Show-o’s original image reconstruction capabilities is minimal, producing reconstructions with significant noise artifacts. Consequently, RecA must simultaneously learn to mitigate noise generation while conditioning on CLIP features, presenting additional optimization challenges.
Why the text-to-image generation improvements on BAGEL is modest? (I) The computational constraints associated with training the 14B parameter model prevent full exploitation of RecA’s potential benefits. (II) We hypothesize that BAGEL’s pre-existing image editing training, which conditions on both SigLIP and VAE features, has already endowed the model with robust capabilities for SigLIP feature-based reconstruction. It constrains the potential improvement space for RecA in generation tasks, while correspondingly enabling substantial enhancements in image editing capabilities.